OpenStudy (anonymous):

A Cold stone Creamery ice cream shop sells sundaes for $3.60 and banana splits for $4.25. The shops sells four times as many sundaes as banana splits. If total sales amounted to $3,730 last weekend, how many banana splits wer sold? What were the total dollar sales of sundaes from previous question?

OpenStudy (anonymous):

200

OpenStudy (nurali):

banana spilit = x sundaes be 4x x(4.25) + 4x(3.60) = 3730 4.25x + 14.40x = 3730 18.65x = 3730 x = 3730/18.65 = 200 banana spilit = 200

OpenStudy (anonymous):

is that sold

OpenStudy (nurali):

Yes

OpenStudy (nurali):

total dollar sales of sundaes is 12